OpenStudy (anonymous):

If f(x) = 5x - 9, find the value of f(-2). A. 8 B. -19 C. 23 D. -23

OpenStudy (haseeb96):

f(-2)=5(-2)-9 =-10-9 =-19 B is the correct answer

OpenStudy (anonymous):

Thanks!